a company recently downsized its number of employees by 30%. If there are still 98 employees, how many employees were there prior to the layoffs.
Answer by Theo(13342)

let x = the number of employees prior to the layoffs.

x - .3 * x = 98
combine like terms to get:
x * (1 - .3) = 98
simplify to get:
.7 * x = 98
divide both sides of the equation by .7 to get:
x = 98 / .7
solve for x to get:
x = 140.

confirm by replacing x with 140 in the original equation to get:
140 - .3 * 140 = 98
simplify to get:
140 - 42 = 98
simplify to get:
98 = 98

this confirms the value of x = 140 is good.

your solution is that the number of employees before the layoff was 140.